OpenStudy (anonymous):

Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2,-2/3)

OpenStudy (anonymous):

@Traxter can you help?

OpenStudy (anonymous):

Sure, the equation has roots 0 and 1, so will look something like this: \[y=ax(x-1)\] For some x which we need to find. Now, you have a corresponding set of x and y values which you can plug into the above equation to find the required a. Are you able to have a go at the next bit yourself or do you need a hand?

OpenStudy (anonymous):

Also are you comfortable with why the equation has the form which I have put it in?

OpenStudy (anonymous):

i need a hand :/

OpenStudy (anonymous):

Ok, substitute x=1/2 and y=-2/3 to get: \[-2/3 = \frac{a}{2}\times(-\frac{1}{2})\] \[\frac{-2}{3}=\frac{-a}{4}\] Multiply by -4 to get: \[a=\frac{8}{3}\] So the equation will be: \[y=\frac{8x}{3}(x-1)\] Do you understand?

OpenStudy (anonymous):

Also, if you expand it will be: \[y=\frac{8x^2}{3}-\frac{8x}{3}\]

OpenStudy (anonymous):

thank you traxter :))

OpenStudy (anonymous):

Not a problem :)